# ReRAM devices, memristive devices, resistance switching memory, optimal control, energy minimization articletype: Paper

### v1

@@ Abstract
Under certain conditions, applying a sequence of voltage pulses of alternating polarities across a resistive switching memory device induces a finite number of fixed-point attractors, known as dynamical attractors. Remarkably, dynamical attractors can be used to program analog values into the device state without supervision. Because different pulse sequences can produce the same trajectory solution for the state in the phase space, there is strong potential for optimization, particularly in regard to the energy cost of the programming phase, which this study addresses. Without loss of generality, the proposed theory-based energy minimization strategy is applied to the voltage threshold adaptive memristor model, known for its predictive capability and adaptability to fit a large number of resistance switching memory devices. The optimization design crafts ad-hoc pulse sequences, that minimize the energy required to program the device into a desired dynamical attractor state. The theoretical approach is also extended to cover situations, where a fast programming scheme should be adopted to serve time-critical electronics applications.
@@
ReRAM devices, memristive devices, resistance switching memory, optimal control, energy minimization
@@ 1 Introduction
The precise programming of resistance switching memory devices [1] is crucial for various applications that depend on memristive states. These applications include analog circuits using memristive devices [2], memristive neural networks [3, 4, 5], and crossbar-based memory arrays that can also function as computing engines, such as accelerating matrix-vector multiplications [6, 7, 8]. On one hand, the effectiveness of memristive technologies heavily relies on the precision during the state programming process. On the other hand, developing methods to reduce power dissipation in a memristive device during programming is critically important, especially in mobile technical systems with limited energy supplies.
In general, applying a sequence of voltage pulses of alternating polarity across a resistance switching memory device can induce a finite number of dynamical attractors in the device’s phase space, opening up novel and intriguing opportunities for device programming. This concept was introduced in [9] and further explored in [10],[11],[12], [13], and [14]. Moreover, in [15], high-frequency periodic signals were used to induce controlled resistance switching transitions in non-volatile memristor devices endowed with fading memory [16]. Importantly, in recent years, several studies explored the theoretical basis behind the synthesis of low-power programming schemes for memristive devices [17, 18, 19]. One notable finding is that the stimulus, minimizing Joule losses in ideal memristors, corresponds to constant power dissipation [17].
@@
This study merges the areas of dynamical attractors and low-power programming of memristive devices. Specifically, we present a theoretical framework that designs a programming pulse sequence to minimize Joule losses within the device while simultaneously producing a desired dynamic attractor. Our analysis is based on the voltage threshold adaptive memristor (VTEAM) model [20] that is a particular realization of first-order memristive devices and systems [21]
@@
This paper is organized as follows. In Sec. 2, we present the VTEAM model and derive a time-averaged trajectory solution for the internal state of a resistance switching memory device, under pulse train excitation. Sec. 3.1 outlines the objectives of the state programming operation, while the remainder of Sec. 3 describes the proposed theoretical methodology for synthesizing the most energy-efficient state programming protocol (Sec. 3.2) and the most energetically-favorable programming scheme under a limited time budget (Sec. 3.3). Joule losses across memristive devices under programming are assessed in Sec. 4. The discussion of our results can be found in Sec. 5. The paper concludes with a conclusion.
@@ 2.1 VTEAM model
where $k_{off}>0$, $k_{on}<0$, $\alpha_{off}>0$, $\alpha_{on}>0$ are constant parameters, whereas $f_{off}(x)$ and $f_{on}(x)$ are window functions, restricting the allowable variation range of the state variable $x$, denoting the normalized width of the insulating part of the functional layer within the device physical stack, to the closed interval $x_{on},x_{off}]$, with $x_{on}=0$ and $x_{off}=1$, and reading as
@@
respectively. According to the state evolution function $f(x,V)$ (Eqs. (3)-(5)), the device undergoes RESET/SET resistance switching transitions when the applied to the device voltage $V$ is above/below a positive/negative threshold level $v_{off}$$v_{on}$.
@@ 2.2 Trajectory
In principle, the Joule losses across the device are defined by the exact trajectory $x(t)$ of the state variable and time-dependent voltage. However, in the limit of narrow pulses, the exact trajectory $x(t)$ may be well approximated by the time-averaged trajectory $\bar{x}(t)$. Under these circumstances the Joule losses can be estimated accurately via
@@ 3.1 Objectives of state programming
Eqs. (18) and (19) demonstrate that to attain greater precision (i.e., smaller $\epsilon$-s), one must employ shorter pulses.
@@ 3.2 The most energy efficient programming scheme
where $V_{-}^{max}$ is the most negative voltage applicable across the device.
@@
Figure 5: Minimizing Joule losses in the presence of constraints on pulse amplitudes and widths in the case where $\alpha_{\text{off}}<2$ and $\alpha_{\text{on}}<2$. The allowable variation ranges for $V_{+}$ and for $V_{-}$ are $(v_{off},V_{+}^{max})$ and $(-v_{on},-V_{-}^{max})$, respectively. (a) A schematic illustration of the scenario where $V_{+}^{max}>2v_{off}/(2-\alpha_{off})$ and $|V_{-}^{max}|<2|v_{on}|/(2-\alpha_{on})$. (b) A schematic illustration of the scenario where $V_{+}^{max}>2\,v_{off}/(2-\alpha_{off})$ and $|V_{-}^{max}|>2|v_{on}|/(2-\alpha_{on})$. It is worth observing that, when both inequalities $\alpha_{\text{on}}<2$ and $\alpha_{\text{off}}<2$ apply, the contours of the function $Q=Q(V_{+},V_{-})$ are closed loci, which originates from the fact that in situations of this kind there always exists finite positive and negative pulse amplitudes, in turn $V_{+}$ and $V_{-}$, at which Joule losses are minimized under RESET and SET transitions, respectively, in the unconstrained case (refer to Fig. 6(b)). On the other hand, in case $\alpha_{\text{on}}$ and/or $\alpha_{\text{off}}$ were larger than or equal to $2$, the contours of the function $Q=Q(V_{+},V_{-})$ would be open curves. In scenarios of this kind, if the first (second) inequality from the aforementioned pair held true, the most energetically-favourable SET (RESET) pulse amplitudes $V_{-}$ ($V_{+}$) would be negative (positive) infinity under unconstrained conditions (refer to Fig. 6(a)). Refer to the text for additional information.
3.3 The most energetically-favourable yet fast programming scheme
@@ 3.3.1 Case 1
Let us now consider the scenario, where Eqs. (22) and (23) result in the optimal pulse amplitude pair choice $\{\hat{V}_{+}=2\,v_{\text{off}}/(2-\alpha_{\text{off}}),\hat{V}_{-}=V_{-}^{max}\}$. This scenario occurs if a) $\alpha_{on}\geq 2$ or $\alpha_{on}<2$ and $|V_{-}^{max}|<2|v_{on}|/(2-\alpha_{on})$ and b) $\alpha_{off}<2$ and $V_{+}^{max}>2\,v_{off}/(2-\alpha_{off})$. Now, let us suppose the memristor-centered application requires the implementation of a fast resistance switching scheme, as dictated by the constraint (26), so as to prevent the programming time to exceed a predefined upper bound. In a situation of this kind, the amplitude for the negative pulse is selected as $\hat{V}_{-}$, $\tau_{-}$ is computed via (19), and, concurrently, the width of the positive pulse is calculated as the remainder of the pulse sequence period, i.e via
@@ 3.3.2 Case 2
$\nabla Q(V_{+},V_{-})=\lambda\nabla\left(\tau_{+}+\tau_{-}\right),$ (30)
where $\nabla$ represents the gradient operator, and $\lambda$ is the Lagrange multiplier. Eq. (30) may be expanded as
@@
The location of the point “S” in Fig. 5(b) can be found by solving simultaneously Eqs. (29) and (31) for $V_{+}$ and $V_{-}$. The widths $\tau_{+}$ and $\tau_{-}$ of the positive and negative pulse are then computed by means of Eqs. (18) and (19), respectively.
@@ 4 Joule losses
Here we point out that Eq. (32) remains invariate across various program protocols (assuming the same trajectory). Variations arise from the pre-integral coefficient in Eq. (13), which must be addressed individually for different pulsing scenarios. In all cases, however, the Joule losses in the memristive device may be accurately evaluated inspecting the closed-form expression
@@
To illustrate the strategy outlined in Sect. 3.2, Fig. 6 displays the programming energy $Q$ versus the pulse amplitude $V_{+}$, relative to the optimal value $Q_{opt}$. This figure was generated for a device with symmetric kinetics, considering each value of the parameters $\alpha_{off}$ of the set $\{0.5,1,1.5\}$ in panel (a) and from the set $\{2,3,4\}$ in panel (b). In the present case, $Q/Q_{opt}$ remains constant regardless of the details of the trajectory, since the $I/T$ factor in Eq. (34) is the same for the same trajectory.
@@
Figure 6: Device energy dissipation $Q$ relative to its optimal counterpart $Q_{opt}$ under positive pulse excitation for $v_{off}=1$ V and $V_{+}^{max}=3$ V, when (a) $\alpha_{off}<2$ and (b) $\alpha_{off}\geq 2$. The calculations were performed for a device with symmetric kinetics (assuming $\alpha_{on}=\alpha_{off}$, $v_{on}=-v_{off}$, $V_{-}^{max}=-V_{-}^{max}$), and the same trajectory $\bar{x}(t)$.
@@ 5 Discussion
Although many experiments on memristive devices have focused on the precision of state programming, another fundamental aspect, namely its energy efficiency, has attracted far less experimental attention [25]. In this study, we have introduced a method to optimize energy cost for programming a non-volatile resistance switching memory device using its dynamical attractor state [9]. This state is reached through the application of a specific sequence of alternating pulses from any initial state, thus in the presence of a fading memory effect, as described in [16]. It has been found that a particular pulses sequence that minimizes the energy cost depends on several model parameters.
@@ 5.2 Joule losses
Figure 6 illustrates the Joule losses predicted by our theoretical model as a function of the amplitude of the RESET pulse. In Fig. 6(a), a scenario is presented where $\alpha_{off}<2$, suggesting a parameter-based RESET pulse configuration for optimal energy efficiency. In this case, the curve $Q/Q_{opt}$ versus $V_{+}$ has a minimum at $V_{+}=2v_{off}/(2-\alpha)$ for any value $\alpha$. However, the minimum’s abscissa of this curve falls within (or outside) the range $(v_{off},V_{+}^{max})$ for $\alpha_{off}\in(0,4/3)$ ($\alpha_{off}\in(4/3,2)$). In the latter situation, $\hat{V}_{+}$ should be set to the upper limit $V_{+}^{max}$ within the permissible range.
@@
The theoretical development of curves like those shown in Fig.  6 can result in substantial energy savings during the critical and essential task of programming resistances in memristive devices.
@@ Conclusion
In conclusion, dynamical attractor states of memristive devices offer an alternative approach to their programming. This research has investigated methods to reduce energy consumption within this framework. Using the VTEAM model, we have evaluated our theory-based optimization strategy to show how energy dissipation in the device can be minimized when exposed to voltage pulses of alternating polarity.
The optimal pulse sequence depends on the critical parameters $\alpha_{off}$ and $\alpha_{on}$, which play a key role in the device’s switching kinetics. When the former (latter) parameter is 2 or higher, the optimal energy-efficient strategy is to select the highest possible magnitude for RESET (SET) pulses. On the other hand, if $\alpha_{off}$ ($\alpha_{on}$) is less than 2, achieving maximum energy efficiency requires setting the RESET (SET) pulse magnitude to the lesser of a certain intermediate voltage magnitude and highest possible voltage magnitude (taking into account the polarity).
Our theory-driven optimization strategy demonstrates that choosing optimized pulse sequences can significantly decrease energy dissipation in memristive devices during repeated programming operations. This theoretical method has also been utilized to explore the reduction of energy dissipation in memristive devices under fast programming schemes, which could be beneficial for real-time edge computing applications. Our theoretical approach is set to be compared with experimental data from real-world implementations of memristive devices in the future.

### v2

# ReRAM devices, memristive devices, resistive switching memory, optimal control, energy minimization articletype: Paper

@@ Abstract
Under certain conditions, applying a sequence of voltage pulses of alternating polarities across a resistive switching memory device induces a finite number of fixed-point attractors in its time-averaged dynamics, known as dynamical attractors. Remarkably, dynamical attractors can be used to program analog values into the device state without supervision. Because different pulse sequences can produce the same trajectory solution for the state in the phase space, there is strong potential for optimization, particularly regarding the energy cost of the programming phase, which this study addresses. The proposed theory-based energy minimization strategy is applied to the voltage threshold adaptive memristor (VTEAM) model, which is known for its predictive capability and adaptability in fitting a large number of resistive switching memory devices. The optimization design crafts ad-hoc pulse sequences that minimize the energy required to program the device into a desired dynamical attractor. The theoretical approach is also extended to cover situations where a fast programming scheme should be adopted to serve time-critical electronics applications.
@@
ReRAM devices, memristive devices, resistive switching memory, optimal control, energy minimization
@@ 1 Introduction
The precise programming of resistive switching memory devices [1] is crucial for various applications that depend on memristive states. These applications include analog circuits using memristive devices [2], memristive neural networks [3, 4, 5], and crossbar-based memory arrays that can also function as computing engines, e.g. for accelerating matrix-vector multiplications [6, 7, 8]. On one hand, the effectiveness of memristive technologies heavily relies on the precision during the state programming process. On the other hand, developing methods to reduce power dissipation in a memristive device during programming is critically important, especially in mobile technical systems with limited energy supplies.
In general, applying a sequence of voltage pulses of alternating polarity across a resistive switching memory device can induce a finite number of dynamical attractors in the device’s phase space, opening up novel and intriguing opportunities for device programming. This concept was introduced in [9] and further explored in [10],[11],[12], and [13]. Moreover, in [14], high-frequency periodic signals were used to induce controlled resistive switching transitions in non-volatile memristor devices endowed with fading memory [15]. Importantly, in recent years, several studies explored the theoretical basis behind the synthesis of low-power programming schemes for memristive devices [16, 17, 18]. One notable finding is that the stimulus, minimizing Joule losses in ideal memristors, corresponds to constant power dissipation [16].
@@
This study merges the areas of dynamical attractors and low-power programming of memristive devices. Specifically, we present a theoretical framework that designs a programming pulse sequence to minimize Joule losses within the device while simultaneously producing a desired dynamic attractor trajectory. Our analysis is based on the voltage threshold adaptive memristor (VTEAM) model [19] that is a particular realization of first-order memristive devices and systems [20]
@@
This paper is organized as follows. In Sec. 2, we present the VTEAM model and derive a time-averaged trajectory solution for the internal state of a resistive switching memory device, under pulse train excitation. Sec. 3.1 outlines the objectives of the state programming operation, while the remainder of Sec. 3 describes the proposed theoretical methodology for synthesizing the most energy-efficient state programming protocol (Sec. 3.2) and the most energetically-favorable programming scheme under a limited time budget (Sec. 3.3). Joule losses across memristive devices under programming are assessed in Sec. 4. The discussion of our results can be found in Sec. 5. The paper concludes with a conclusion.
@@ 2.1 VTEAM model
where $k_{off}>0$, $k_{on}<0$, $\alpha_{off}>0$, $\alpha_{on}>0$ are constant parameters, whereas $f_{off}(x)$ and $f_{on}(x)$ are window functions, restricting the allowable variation range of the state variable $x$, that, in some physical devices, represents the normalized distance between the tip of a conductive filament and the electrode opposite to the one, the filament itself originated from, to the closed interval $x_{on},x_{off}]$, with $x_{on}=0$ and $x_{off}=1$, and reading as
@@
respectively. According to the state evolution function $f(x,V)$ (Eqs. (3)-(5)), the device undergoes RESET/SET resistive switching transitions when the voltage V, applied across the device, is above/below a positive/negative threshold level $v_{off}$$v_{on}$.
@@
We note that VTEAM [19] is a phenomenological model whose parameters are obtained by fitting measurement results. In particular, Ref. [19] reports parameters for several experimental resistive switching memory devices.
@@ 2.2 Trajectory
In this work, we focus on programming devices with square pulses (as illustrated in Fig. 1(a)), although other pulse shapes can also be used for this purpose. Square pulses are experimentally simpler to implement, more straightforward to analyze, and, when sufficient programming time is available, they are always optimal for switching [18].
@@
In principle, the Joule losses across the device are defined by the exact trajectory $x(t)$ of the state variable and the applied time-dependent voltage $V(t)$. However, in the limit of narrow pulses, the exact trajectory $x(t)$ may be well approximated by the time-averaged trajectory $\bar{x}(t)$. Under these circumstances the Joule losses can be estimated accurately via
@@ 3.1 Objectives of state programming
Eqs. (18) and (19) demonstrate that to attain greater precision (i.e., smaller $\epsilon$), one must employ shorter pulses.
@@ 3.2 The most energy efficient programming scheme
where $V_{-}^{max}$ is the most negative voltage applicable across the device. $V_{+/-}^{max}$ should be selected taking into account the largest voltage or voltages available in the circuit and the operational voltage stress thresholds for the device.
@@
Figure 5: Minimizing Joule losses in the presence of constraints on pulse amplitudes and widths in the case where $\alpha_{\text{off}}<2$ and $\alpha_{\text{on}}<2$. The allowable variation ranges for $V_{+}$ and for $V_{-}$ are $v_{off},V_{+}^{max}]$ and $v_{on},-V_{-}^{max}]$, respectively. (a) A schematic illustration of the scenario where $V_{+}^{max}>2v_{off}/(2-\alpha_{off})$ and $|V_{-}^{max}|<2|v_{on}|/(2-\alpha_{on})$. (b) A schematic illustration of the scenario where $V_{+}^{max}>2\,v_{off}/(2-\alpha_{off})$ and $|V_{-}^{max}|>2|v_{on}|/(2-\alpha_{on})$. It is worth observing that, when both inequalities $\alpha_{\text{on}}<2$ and $\alpha_{\text{off}}<2$ apply, the contours of the function $Q=Q(V_{+},V_{-})$ are closed loci, which originates from the fact that in situations of this kind there always exist finite positive and negative pulse amplitudes, in turn $V_{+}$ and $V_{-}$, at which Joule losses are minimized under RESET and SET transitions, respectively, in the unconstrained case (refer to Fig. 6(a)).
3.3 The most energy efficient programming scheme under a limited time budget
@@ 3.3.1 Case 1
Let us now consider the scenario, where Eqs. (22) and (23) result in the optimal pulse amplitude pair choice $\{\hat{V}_{+}=2\,v_{\text{off}}/(2-\alpha_{\text{off}}),\hat{V}_{-}=V_{-}^{max}\}$. This scenario occurs if a) $\alpha_{on}\geq 2$ or $\alpha_{on}<2$ and $|V_{-}^{max}|<2|v_{on}|/(2-\alpha_{on})$ and b) $\alpha_{off}<2$ and $V_{+}^{max}>2\,v_{off}/(2-\alpha_{off})$. Now, let us suppose the memristor-centered application requires the implementation of a fast resistive switching scheme, as dictated by the constraint (26), so as to prevent the programming time to exceed a predefined upper bound. In a situation of this kind, the amplitude for the negative pulse is selected as $\hat{V}_{-}$, $\tau_{-}$ is computed via (19), and, concurrently, the width of the positive pulse is calculated as the remainder of the pulse sequence period, i.e via
@@ 3.3.2 Case 2
$\vec{\nabla}Q(V_{+},V_{-})=\lambda\vec{\nabla}\left(\tau_{+}+\tau_{-}\right),$ (30)
where $\vec{\nabla}$ represents the gradient operator, and $\lambda$ is the Lagrange multiplier. Eq. (30) may be expanded as
@@
The location of the point “S” in Fig. 5(b) can be found by solving simultaneously Eqs. (29) and (31) for $V_{+}$ and $V_{-}$. The widths $\tau_{+}$ and $\tau_{-}$ of the positive and negative pulses are then computed by means of Eqs. (18) and (19), respectively.
@@ 4 Joule losses
Here we point out that Eq. (32) remains invariant across various programming protocols (assuming the same trajectory). Variations arise from the pre-integral coefficient in Eq. (13), which must be addressed individually for different pulse sequences. In all cases, however, the Joule losses in the memristive device may be accurately evaluated inspecting the closed-form expression
@@
To illustrate the strategy outlined in Sect. 3.2, Fig. 6 displays the programming energy $Q$ versus the pulse amplitude $V_{+}$, relative to the optimal value $Q_{opt}$. This figure was generated for a device with symmetric kinetics, considering each value of the parameters $\alpha_{off}$ from the set $\{0.5,1,1.5\}$ in panel (a) and from the set $\{2,3,4\}$ in panel (b). In the present case, $Q/Q_{opt}$ remains constant regardless of the details of the trajectory, since the $I/T$ factor in Eq. (34) is the same for the same trajectory $\bar{x}(t)$.
@@
Figure 6: Energy dissipation relative to the optimal dissipation found for (a) $\alpha_{off/on}<2$ and (b) $\alpha_{off/on}\geq 2$. These graphs were obtained using $v_{off}=1$ V, $v_{on}=-1$ V, $\alpha_{on}=\alpha_{off}$, $_{+}^{max}=3 V, V$$_{-}^{max}=-3 V$, and $V_{-}=-V_{+}$.
@@ 5 Discussion
Although many experiments on memristive devices have focused on the precision of state programming, another fundamental aspect, namely its energy efficiency, has attracted far less experimental attention [24]. In this study, we have introduced a method to optimize the energy cost for programming a non-volatile resistive switching memory device using its dynamical attractor state [9]. This state is reached through the application of a specific sequence of alternating pulses from any initial state, thus in the presence of a fading memory effect, as described in [15]. It has been found that the particular pulse sequence, that minimizes the programming energy cost, depends upon several model parameters.
@@ 5.2 Joule losses
Figure 6 illustrates the Joule losses predicted by our theoretical model as a function of the amplitude of the RESET pulse. In Fig. 6(a), a scenario is presented where $\alpha_{off}<2$, suggesting a parameter-based RESET pulse configuration for optimal energy efficiency. In this case, the curve $Q/Q_{opt}$ versus $V_{+}$ has a minimum at $V_{+}=2v_{off}/(2-\alpha)$ for any value $\alpha$. However, the abscissa of the minimum of this curve falls within (outside) the range $(v_{off},V_{+}^{max})$ for $\alpha_{off}\in(0,4/3)$ ($\alpha_{off}\in(4/3,2)$). In the latter situation, $\hat{V}_{+}$ should be set to the upper limit $V_{+}^{max}$ within the permissible range.
@@
Taking into account the properties of the curves in Fig. 6 can result in substantial energy savings during the critical and essential task of programming resistances in memristive devices.
5.3 Endurance and device-to-device variability effects
This study assumes that the endurance is sufficient to perform programming via the dynamical attractor states throughout the device lifetime. In fact, a sequence of short pulses separated by zero voltage intervals is expected to cause less damage to the device than a single long pulse that produces the same change in $x$, because the long pulse deposits energy within a shorter time window. From a theoretical standpoint, however, determining the effects of device endurance limitations on the proposed dynamic attractor state-based programming scheme is a difficult task. A study of the impact of device-to-device variability on the accuracy of our programming paradigm can be found in the Supplementary Information file.
@@ Conclusion
In conclusion, the dynamical attractor states of memristive devices offer an alternative unsupervised approach to their programming. This research has investigated methods to reduce energy consumption within this framework focusing on devices described by the VTEAM model. Our approach can be extended to other device models and situations, including programming using current pulses. Moreover, a range of practical device limitations—such as variability, zero-input resistance drift, and environmental influences—can be incorporated into future modeling of dynamical attractors.
The optimal pulse sequence depends on the critical parameters $\alpha_{off}$ and $\alpha_{on}$ of the model, which play a key role in the device’s switching kinetics. When the former (latter) parameter is 2 or higher, the optimal energy-efficient strategy envisages the selection of the highest possible magnitude for RESET (SET) pulses (fast switching). On the other hand, if $\alpha_{off}$ ($\alpha_{on}$) is less than 2, achieving maximum energy efficiency requires setting the RESET (SET) pulse magnitude to the lesser of a certain intermediate voltage magnitude (slower switching) and the highest possible voltage magnitude (taking into account the polarity).
The results of this study demonstrate that the choice of optimized pulse sequences can significantly decrease energy dissipation in memristive devices during programming using their dynamical attractors. This theoretical method has also been utilized to explore the reduction of energy dissipation in memristive devices under fast programming schemes, which could be beneficial for real-time edge computing applications. We hope our theoretical findings will motivate future experimental investigations aimed at evaluating the benefits of the proposed approach to program a non-volatile memristor device.